Equipment was acquired at the beginning of the year at a cost of $76,020. The equipment was depreciated using the straight-line method based upon an estimated useful life of 6 years and an estimated residual value of $7,800. 


Required:

 a. What was the depreciation expense for the first year?

 b. Assuming the equipment was sold at the end of the second year for $57,542, determine the gain or loss on sale of the equipment.

 c. Joumalize the entry to record the sale. Refer to the Chart of Accounts for exact wording of account titles. 
Answer #1

a) Depreciation for the First Year= (Equipment Value-Residual Value) / No.of Years

= \frac{76020-7800}{6} = $11370

b) Accumulated Depreciation for the second year end= 11370 X 2= $22740

  Value of Ewuipment at the end of Second Year= 76020-22740= $53280

Sale Price of Equipment= 57542

Gain= 57542-53280= $4262

c)

Cash a/c 57542
Accumulated Depreciation a/c 22740
Gain on sale of Equipment a/c 4262
Equipment a/c 76020
Total 80282 80282